At what speed will an object's kinetic energy be 23 % of its rest energy?
Answer #1

rest energy is given by

E = mc^2

now according to question

(1/2)mv^2 = (23/100)*mc^2

v^2 = (46/100)c^2

v = 0.678*c

v = 0.678*3*10^8 = 20.34*10^7 m/s